# Quantum density estimation with density matrices: Application to quantum anomaly detection

### 22 Dec 2023 v4

@@ VI.2 QAD: Results and discussion
The results obtained by applying the proposed methods for hybrid anomaly detection with the Pennylane quantum simulator are reported in Tables 7 and 8. Compared to QRFF, the QAFF obtained better results in the three evaluated metrics, F1 Score, accuracy, and AUC, thanks to the optimization of the Fourier weights. Regarding the standard deviations, we observe more consistent results across different experiments with QAFF, which shows that QRFF is more sensitive to randomness than QAFF in a low-dimensional setting.
@@ 4 QAFF
Table 7: Obtained metrics for 10 hybrid anomaly detection experiments on Cardio data set using Q-DEMDE method with 4 Fourier features for both QRFF and QAFF with the Pennylane quantum simulator.
@@ 8 QRFF
Q-DEMDE $\mathbf{0.573\pm 0.083}$ $\mathbf{0.920\pm 0.010}$ $\mathbf{0.915\pm 0.028}$
@@ 8 QAFF
Table 8: Obtained metrics for 10 hybrid anomaly detection experiments on Cardio data set using Q-DEMDE method with 8 Fourier features for both QRFF and QAFF with the Pennylane quantum simulator.
Furthermore, in Table 9, we present the results of the hybrid anomaly detection algorithm on the real IBM-Oslo quantum computer; we report only one experiment for each quantum mapping. The results follow a similar pattern as in the quantum simulator: QAFF leads to higher performance than QRFF. In contrast to the quantum simulator, the experiment with QRFF on the real quantum computer presented inferior results, probably because the density predictions with random features had low variance, making it difficult to make a distinction between the two classes due to quantum noise. Nevertheless, the Q-DEMDE with QAFF in the real quantum computer obtained satisfactory results for the anomaly detection task, showing that this strategy is suitable for current quantum devices.
@@
Table 9: Obtained metrics for the hybrid anomaly detection experiments using a real IBM quantum computer (IBM-Oslo) using the Q-DEMDE density estimation strategy with both QRFF and QAFF with 4 Fourier components. We used a classical computer to learn the optimal QAFF, and the real quantum computer to predict the density of validation and test samples with the Q-DMKDE quantum circuit.
@@ VII Conclusions
In this article, we presented a quantum-classical density estimation method called Q-DEMDE for current quantum computers based on the expected values of density matrices and quantum Fourier features. Within our method, we proposed a quantum protocol for mixed-state preparation and expected value estimation, which was used as a quantum subroutine to implement in quantum hardware the classical density estimation algorithm DMKDE [27]. We also introduced two quantum feature maps called quantum random and quantum adaptive Fourier features, which allow to approximate a Gaussian kernel between data samples in both classical and quantum computers. We performed several experiments on various density estimation tasks on a quantum simulator and a real quantum computer.
Experimental results show that it is possible to exploit the properties of mixed states and Fourier features to efficiently represent probability distributions of data in quantum hardware. A key feature of the proposed hybrid DE method is that it is non-parametric, i.e., it does not assume a prior probability distribution to fit the training data, in contrast to parametric quantum DE methods [17, 18]. Furthermore, the Q-DEMDE method with the QRFF map is optimization-free, i.e., it does not require any optimization, while the Q-DEMDE method with QAFF requires optimization only to construct the quantum map. Although the Q-DEMDE algorithm does not achieve a quantum advantage over the classical DMKDE, it shows that it is possible to perform quantum-classical density estimation with a reduced number of quantum Fourier features and relatively small quantum circuits, illustrating its viability for present-day quantum computers.
@@ Acknowledgments
We acknowledge the use of IBM Quantum services for this work. The views expressed in this work are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team.
@@ Appendix A Quantum adaptive Fourier features in a quantum computer with one and two qubit gates
We illustrate the proposed method starting from base cases starting with the method to prepare the state |$\ket{\psi(\bm{x})}_{2}=\sqrt{\frac{1}{4}}\begin{bmatrix}1,e^{i\sqrt{\gamma}\mathbf{w}_{1}\cdot\bm{x}},e^{i\sqrt{\gamma}\mathbf{w}_{2}\cdot\bm{x}},e^{i\sqrt{\gamma}\mathbf{w}_{3}\cdot\bm{x}}\end{bmatrix}$ that corresponds to 2 qubits, 4 QAFF and the classical feature $\bm{x}\in\mathbb{R}^{D}$. The method starts by applying a single Hadamard gate to each qubit, followed by rotations $R_{z}(\sqrt{\gamma}\bm{\theta}_{1}\cdot\bm{x})$ in the first qubit and $R_{z}(\sqrt{\gamma}\bm{\theta}_{2}\cdot\bm{x})$ in the second qubit. We then apply a CNOT gate with control the first qubit and target the second qubit followed by a last rotation $R_{z}(\sqrt{\gamma}\bm{\theta}_{3}\cdot\bm{x})$ on the second qubit as show in the Fig. 9a. These operations create the state |$\ket{\psi(\bm{x})}_{2}=\sqrt{\frac{1}{4}}\begin{bmatrix}1,e^{i\sqrt{\gamma}(\bm{\theta}_{1}+\bm{\theta}_{2})\cdot\bm{x}},e^{i\sqrt{\gamma}(\bm{\theta}_{2}+\bm{\theta}_{3})\cdot\bm{x}},e^{i\sqrt{\gamma}(\bm{\theta}_{1}+\bm{\theta}_{3})\cdot\bm{x}}\end{bmatrix}^{\prime}$. The structure of this quantum state preparation for two qubits has been explored in Ref. [74] in the context of amplitude encoding. The desired quantum Fourier map is obtained by setting the parameters $\{\bm{\theta}_{j}\}$ such that $\bm{\theta}_{1}+\bm{\theta}_{2}=\mathbf{w}_{1}$, $\bm{\theta}_{2}+\bm{\theta}_{3}=\mathbf{w}_{2}$, and $\bm{\theta}_{1}+\bm{\theta}_{3}=\mathbf{w}_{3}$.
@@
Figure 9: Proposed quantum circuits for the Quantum Adaptive Fourier Features operation. a) shows the circuit for two qubits and a size of 4 QAFF, and b) shows the circuit for three qubits and 8 QAFF. As an example of the recursive construction of these operations, c) show the operation $C_{4}^{1,2,3\rightarrow 4}$ of the circuit for 16 QAFF, formed by two copies of the block $C_{3}^{1,2\rightarrow 3}$ in b) and a CNOT gate.
@@
It is also possible to rewrite the previous state preparation by defining the operations $C_{1}^{0\rightarrow 1}$ and $C^{1\rightarrow 2}_{2}$ as controlled rotations in the $z$ axis, where the subindex indicates the number of qubits involved in the operation, and the superindex indicates the control qubits followed by an arrow to the single target qubit, with the top qubit of the circuit considered as the qubit 1, and the cero index indicating no controlled qubit. The two-qubit state preparation can be realized by applying a $H$ gate to each qubit followed by the $C_{1}^{0\rightarrow 1}$ operation in the first qubit corresponding to a $R_{z}$ rotation, followed by the $C^{1\rightarrow 2}_{2}$, which is built by grouping the two single $R_{z}$ rotations in the second qubit with the CNOT, as shown in Fig. 9a.
@@
We now illustrate the quantum state preparation for eight quantum Fourier features and three qubits with single $R_{z}$ qubit rotations with angles $\{\sqrt{\gamma}\bm{\theta}_{j}\cdot\bm{x}\}_{j=1,\cdots,7}$. First, we apply a Hadamard gate to each of the three qubits, and then the preparation of the first two qubits corresponds to the previous two-qubit state preparation, i.e., we apply the operations $C_{1}^{0\rightarrow 1}$, and $C^{1\rightarrow 2}_{2}$, and then we apply twice the operation $C^{1\rightarrow 3}_{2}$ that shares the same structure of $C^{1\rightarrow 2}_{2}$; the CNOT has the same control qubit (first qubit), but it changes the $R_{z}$ rotations and the target of the CNOT from qubit 2 to qubit 3. Hence, we apply twice $C^{1\rightarrow 3}_{2}$ with different angular rotations in the $z$ basis as shown in Fig. 9b and then we connect these operations with a CNOT with control the 2nd qubit and target the 3rd qubit, creating the state,
@@
We may generalize these previous state preparation strategies by defining recursively the operation $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$, which corresponds to a transformation on $n$ qubits with controls the first $n-1$ qubits and target the $n^{\text{th}}$ qubit. This operation is defined by considering the previous case $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n-1}}$ whose target is the $(n-1)^{\text{th}}$ qubit. We may construct an equivalent operation $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n}}$ that shares the same sequence of gates as the $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n-1}}$, namely, both share the same control qubits for the CNOTs, but the $R_{z}$ rotations and targets of the CNOTs move from the $(n-1)^{\text{th}}$ qubit to the $n^{\text{th}}$ qubit. Then, we construct recursively $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$, by applying twice $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n}}$ on the $n$ qubits, each duplicate with distinct $R_{z}$ rotation angles and connecting these two duplicates with a CNOT with control $(n-1)^{\text{th}}$ qubit and target the $n^{\text{th}}$ qubit as shown in Fig. 10a. To further illustrate the method, Fig. 9c shows the $C^{1,2,3\rightarrow 4}_{4}$ case which is defined from $C^{1,2\rightarrow 3}_{3}$ (Fig. 9b).
Figure 10: Recursive construction of the quantum adaptive Fourier features on $n$ qubits. a) shows the construction of the operator $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$ as the concatenation of two operators $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow n}$ and a CNOT gate in between. b) shows the construction of the total QAFF circuit for an arbitrary number of qubits $n$ in terms of these recursive operators.
Based on the transformations $\{C_{i}^{{1,\cdots,{i-1}}\rightarrow i}\}$, we may construct an ansatz for QAFF that also corresponds to an arbitrary quantum state preparation in the $z$ basis on $n$ qubits, by first applying a $H$ gate to each of the $n$ qubits followed by the $C_{1}^{1}$, $C^{1\rightarrow 2}_{2}$, $C^{1,2\rightarrow 3}_{3}$, $C^{1,2,3\rightarrow 4}_{4}$, $\cdots$, $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$. The final quantum state |$\ket{\psi(\bm{x})}_{n}$ will be given by,
@@
As stated before, this recursive method can also be used to build an ansatz for an arbitrary pure quantum state on $n$ qubits. To build an arbitrary state in the $z$ basis, we apply the same ansatz as the QAFF (which starts by applying $H_{n}$ to the |$\ket{0}_{n}$ state), but replace the single-qubit rotations $\{R_{z}(\sqrt{\gamma}{\bm{\theta}_{j}}\cdot\bm{x})\}_{j=1,\cdots d-1}$ with the rotations $\{R_{z}(\delta_{j})\}_{j=1,\cdots d-1}$, where the $\{\delta_{j}\}$ depend on the phases of the pure state to be learned or to prepared and $d=2^{n}$. We can also build a pure quantum state with only real components by ignoring the $H_{n}$ operation, and applying the sequence of $\{C_{i}^{{1,\cdots,{i-1}}\rightarrow i}\}$ to the |$\ket{0}_{n}$ state with $R_{y}$ rotations instead of $R_{z}$ rotations. Furthermore, the QAFF ansatz can be used to prepare an arbitrary complex-valued quantum state of $n$ qubits by first building a pure state with real components with the $R_{y}$ rotations followed by the $R_{z}$ rotations to account for the phases of the complex components. Compared to a common mechanism for building quantum states such as the presented in Ref. [57], our method uses fewer CNOT gates for the preparation of arbitrary quantum states in either the $y$ basis or in the $z$ basis, but has the same depth and structure when combining both $y$ and $z$ rotations to build arbitrary quantum states.
Considering that the depth of the QAFF map |$\ket{\psi(\bm{x})}_{n}$ (number of CNOTs) is $2^{n}-(n+1)=d-(\log{d}-1)$ ), which is independent on the size $D$ of the original data, this mapping can represent high-dimensional data in quantum computers with a low number of qubits. Furthermore, the complexity of the QAFF method for estimating the Gaussian kernel with the projection |$\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ is $O(TDd+TdR)$ where $T$ is the number of data pairs $\{(\bm{x}_{l},\bm{x}_{m})\}$, and $R$ is the total number of |$\ket{0}_{n}$ measurements required to estimate $\hat{k}_{\{\mathbf{w}_{j}\}}(\bm{x}_{l},\bm{x}_{m})=\absolutevalue{\innerproduct{\psi(\bm{x}_{l})}{\psi(\bm{x}_{l})}{\psi(\bm{x}_{m})}}^{2}$ with the quantum kernel estimation algorithm [59]; this result follows from the fact that for each data pair $(\bm{x}_{l},\bm{x}_{m})$ it is required $O(Dd)$ steps to find the angles $\{\sqrt{\gamma}\bm{\theta}_{j}\cdot\bm{x}\}$, $O(d)$ to prepare the QAFF states and $R$ measurement shots to estimate the projection. In Appendix B, we illustrate the overall quantum complexity of training QAFF by taking into account the gradients of the method.

### 18 Mar 2024 v5

@@ VI.2 QAD: Results and discussion
The results obtained by applying the proposed methods for hybrid anomaly detection with the Pennylane quantum simulator are reported in Tables 7 and 8. Compared to QRFF, the QAFF obtained better results in the three evaluated metrics, F1 Score, accuracy, and AUC, thanks to the optimization of the Fourier weights. Regarding the standard deviations, we observe more consistent results across different trials with QAFF, which shows that QRFF is more sensitive to randomness than QAFF in a low-dimensional setting.
@@ 4 QAFF
Table 7: Obtained metrics for 10 hybrid anomaly detection trials on Cardio data set using Q-DEMDE method with 4 Fourier features for both QRFF and QAFF with the Pennylane quantum simulator.
@@ 8 QRFF
Q-DEMDE $\mathbf{0.573\pm 0.083}$ $\mathbf{0.920\pm 0.010}$ $0.915\pm 0.028$
@@ 8 QAFF
Table 8: Obtained metrics for 10 hybrid anomaly detection trials on Cardio data set using Q-DEMDE method with 8 Fourier features for both QRFF and QAFF with the Pennylane quantum simulator.
Furthermore, in Table 9, we present the results of the hybrid anomaly detection algorithm on the real IBM-Oslo quantum computer; we performed only one demonstration of the QAD strategy for each quantum mapping. The results follow a similar pattern as in the quantum simulator: QAFF leads to higher performance than QRFF. In contrast to the quantum simulator, the demonstration with QRFF on the real quantum computer presented inferior results, probably because the density predictions with random features had low variance, making it difficult to make a distinction between the two classes due to quantum noise. Nevertheless, the Q-DEMDE with QAFF in the real quantum computer obtained satisfactory results for the anomaly detection task, showing that this strategy is suitable for current quantum devices.
@@
Table 9: Obtained metrics for hybrid anomaly detection on the real IBM-Oslo quantum computer using the Q-DEMDE density estimation strategy with both QRFF and QAFF with 4 Fourier components. We used a classical computer to learn the optimal QAFF, and the real quantum computer to predict the density of validation and test samples with the Q-DMKDE quantum circuit.
@@ VII Conclusions
In this article, we presented a quantum-classical density estimation method called Q-DEMDE for current quantum computers based on the expected values of density matrices and quantum Fourier features. Within our method, we proposed a quantum protocol for mixed-state preparation and expected value estimation, which was used as a subroutine to implement the classical density estimation algorithm DMKDE [27] in quantum hardware. We also introduced two quantum feature maps called quantum random and quantum adaptive Fourier features, which allow to approximate a Gaussian kernel between data samples in both classical and quantum computers. We performed various demonstrations on different data sets for density estimation on a quantum simulator and on a real quantum computer.
The quantum-classical density estimation results show that it is possible to exploit the properties of mixed states and Fourier features to efficiently represent probability distributions of data in quantum hardware. A key feature of the proposed hybrid DE method is that it is non-parametric, i.e., it does not assume a prior probability distribution to fit the training data, in contrast to parametric quantum DE methods [17, 18]. Furthermore, the Q-DEMDE method with the QRFF map is optimization-free, i.e., it does not require any optimization, while the Q-DEMDE method with QAFF requires optimization only to construct the quantum map. Although the Q-DEMDE algorithm does not achieve a quantum advantage over the classical DMKDE, it shows that it is possible to perform quantum-classical density estimation with a reduced number of quantum Fourier features and relatively small quantum circuits, illustrating its viability for present-day quantum computers.
@@
Code availability
The source code necessary to replicate the demonstrations presented in this paper is available in the following GitHub repository https://github.com/diegour1/QDEMDE.
@@ Acknowledgments
We acknowledge the use of IBM Quantum services for this work. The views expressed in this work are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum team. D.H.U. acknowledges Google Research for the support through the Google Ph.D. Fellowship.
@@ Appendix A Quantum adaptive Fourier features in a quantum computer with one and two qubit gates
We illustrate the proposed method starting from base cases starting with the method to prepare the state |$\ket{\psi(\bm{x})}_{2}=\sqrt{\frac{1}{4}}\begin{bmatrix}1,e^{i\sqrt{\gamma}\mathbf{w}_{1}\cdot\bm{x}},e^{i\sqrt{\gamma}\mathbf{w}_{2}\cdot\bm{x}},e^{i\sqrt{\gamma}\mathbf{w}_{3}\cdot\bm{x}}\end{bmatrix}$ that corresponds to 2 qubits, 4 QAFF and the classical feature $\bm{x}\in\mathbb{R}^{D}$. The method starts by applying a single Hadamard gate to each qubit, followed by rotations $R_{z}(\sqrt{\gamma}\bm{\theta}_{1}\cdot\bm{x})$ in the first qubit and $R_{z}(\sqrt{\gamma}\bm{\theta}_{2}\cdot\bm{x})$ in the second qubit. We then apply a CNOT gate which controls the first qubit and targets the second qubit followed by a last rotation $R_{z}(\sqrt{\gamma}\bm{\theta}_{3}\cdot\bm{x})$ on the second qubit as show in the Fig. 9a. These operations create the state |$\ket{\psi(\bm{x})}_{2}=\sqrt{\frac{1}{4}}\begin{bmatrix}1,e^{i\sqrt{\gamma}(\bm{\theta}_{1}+\bm{\theta}_{2})\cdot\bm{x}},e^{i\sqrt{\gamma}(\bm{\theta}_{2}+\bm{\theta}_{3})\cdot\bm{x}},e^{i\sqrt{\gamma}(\bm{\theta}_{1}+\bm{\theta}_{3})\cdot\bm{x}}\end{bmatrix}^{\prime}$. The structure of this quantum state preparation for two qubits has been explored in Ref. [75] in the context of amplitude encoding. The desired quantum Fourier map is obtained by setting the parameters $\{\bm{\theta}_{j}\}$ such that $\bm{\theta}_{1}+\bm{\theta}_{2}=\mathbf{w}_{1}$, $\bm{\theta}_{2}+\bm{\theta}_{3}=\mathbf{w}_{2}$, and $\bm{\theta}_{1}+\bm{\theta}_{3}=\mathbf{w}_{3}$.
@@
Figure 9: Proposed quantum circuits for the Quantum adaptive Fourier features operation, (a) shows the circuit for two qubits and a size of 4 QAFF, and (b) shows the circuit for three qubits and 8 QAFF. As an example of the recursive construction of these operations, (c) illustrates the operation $C_{4}^{1,2,3\rightarrow 4}$ of the circuit for 16 QAFF, formed by two copies of the block $C_{3}^{1,2\rightarrow 3}$ in (b) and a CNOT gate.
@@
It is also possible to rewrite the previous state preparation by defining the operations $C_{1}^{0\rightarrow 1}$ and $C^{1\rightarrow 2}_{2}$ as controlled rotations in the $z$ axis, where the subindex indicates the number of qubits involved in the operation, and the superindex indicates the controlled qubits followed by an arrow to the single target qubit, with the top qubit of the circuit considered as the qubit 1, and the cero index indicating no controlled qubit. The two-qubit state preparation can be realized by applying a $H$ gate to each qubit followed by the $C_{1}^{0\rightarrow 1}$ operation in the first qubit corresponding to a $R_{z}$ rotation, followed by the $C^{1\rightarrow 2}_{2}$, which is built by grouping the two single $R_{z}$ rotations in the second qubit with the CNOT, as shown in Fig. 9a.
@@
We now illustrate the quantum state preparation for eight quantum Fourier features and three qubits with single $R_{z}$ qubit rotations with angles $\{\sqrt{\gamma}\bm{\theta}_{j}\cdot\bm{x}\}_{j=1,\cdots,7}$. First, we apply a Hadamard gate to each of the three qubits, and then the preparation of the first two qubits corresponds to the previous two-qubit state preparation, i.e., we apply the operations $C_{1}^{0\rightarrow 1}$, and $C^{1\rightarrow 2}_{2}$, and then we apply twice the operation $C^{1\rightarrow 3}_{2}$ that shares the same structure of $C^{1\rightarrow 2}_{2}$; the CNOT has the same control qubit (first qubit), but it changes the $R_{z}$ rotations and the target of the CNOT from qubit 2 to qubit 3. Hence, we apply twice $C^{1\rightarrow 3}_{2}$ with different angular rotations in the $z$ basis as shown in Fig. 9b and then we connect these operations with a CNOT which controls the second qubit and targets the third qubit, creating the state,
@@
We may generalize these previous state preparation strategies by defining recursively the operation $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$, which corresponds to a transformation on $n$ qubits with controls the first $n-1$ qubits and target the $n^{\text{th}}$ qubit. This operation is defined by considering the previous case $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n-1}}$ whose target is the $(n-1)^{\text{th}}$ qubit. We may construct an equivalent operation $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n}}$ that shares the same sequence of gates as the $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n-1}}$, namely, both share the same control qubits for the CNOTs, but the $R_{z}$ rotations and targets of the CNOTs move from the $(n-1)^{\text{th}}$ qubit to the $n^{\text{th}}$ qubit. Then, we construct recursively $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$, by applying twice $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow{n}}$ on the $n$ qubits, each duplicate with distinct $R_{z}$ rotation angles and connecting these two duplicates with a CNOT which controls the $(n-1)$th qubit and targets the $n$th qubit as shown in Fig. 10a. To further illustrate the method, Fig. 9c shows the $C^{1,2,3\rightarrow 4}_{4}$ case which is defined from $C^{1,2\rightarrow 3}_{3}$ (Fig. 9b).
Figure 10: Recursive construction of the quantum adaptive Fourier features on $n$ qubits, (a) shows the construction of the operator $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$ as the concatenation of two operators $C_{n-1}^{{1,\cdots,{n-2}}\rightarrow n}$ and a CNOT gate in between which controls the $(n-1)$th qubit and targets the $(n)$th qubit, and (b) shows the construction of the total QAFF circuit for an arbitrary number of qubits $n$ in terms of these recursive operators.
Based on the transformations $\{C_{i}^{{1,\cdots,{i-1}}\rightarrow i}\}$, we may construct an ansatz for QAFF that also corresponds to an arbitrary quantum state preparation in the $z$ basis on $n$ qubits, by first applying a $H$ gate to each of the $n$ qubits followed by the $C_{1}^{0\rightarrow 1}$, $C^{1\rightarrow 2}_{2}$, $C^{1,2\rightarrow 3}_{3}$, $C^{1,2,3\rightarrow 4}_{4}$, $\cdots$, $C_{n}^{{1,\cdots,{n-1}}\rightarrow n}$. The final quantum state |$\ket{\psi(\bm{x})}_{n}$ will be given by,
@@
As stated before, this recursive method can also be used to build an ansatz for an arbitrary pure quantum state on $n$ qubits. To build an arbitrary state in the $z$ basis, we apply the same ansatz as the QAFF (which starts by applying $H_{n}$ to the |$\ket{0}_{n}$ state), but replace the single-qubit rotations $\{R_{z}(\sqrt{\gamma}{\bm{\theta}_{j}}\cdot\bm{x})\}_{j=1,\cdots d-1}$ with the rotations $\{R_{z}(\delta_{j})\}_{j=1,\cdots d-1}$, where the $\{\delta_{j}\}$ depend on the phases of the pure state to be learned or to be prepared and $d=2^{n}$. We can also build a pure quantum state with only real components by ignoring the $H_{n}$ operation, and applying the sequence of $\{C_{i}^{{1,\cdots,{i-1}}\rightarrow i}\}$ to the |$\ket{0}_{n}$ state with $R_{y}$ rotations instead of $R_{z}$ rotations. Furthermore, the QAFF ansatz can be used to prepare an arbitrary complex-valued quantum state of $n$ qubits by first building a pure state with real components with the $R_{y}$ rotations followed by the $R_{z}$ rotations to account for the phases of the complex components. Compared to a common mechanism for building quantum states such as the presented in Ref. [58], our method uses fewer CNOT gates for the preparation of arbitrary quantum states in either the $y$ basis or in the $z$ basis, but has the same depth and structure when combining both $y$ and $z$ rotations to build arbitrary quantum states.
Considering that the depth of the QAFF map |$\ket{\psi(\bm{x})}_{n}$ (number of CNOTs) is $2^{n}-(n+1)=d-(\log{d}-1)$ ), which is independent on the size $D$ of the original data, this mapping can represent high-dimensional data in quantum computers with a low number of qubits. Furthermore, the complexity of the QAFF method for estimating the Gaussian kernel with the projection |$\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ is $O(TDd+TdR)$ where $T$ is the number of data pairs $\{(\bm{x}_{l},\bm{x}_{m})\}$, and $R$ is the total number of |$\ket{0}_{n}$ measurements required to estimate $\hat{k}_{\{\mathbf{w}_{j}\}}(\bm{x}_{l},\bm{x}_{m})=\absolutevalue{\langle\psi(\bm{x}_{l})|\psi(\bm{x}_{m})\rangle}^{2}$ with the quantum kernel estimation algorithm [60]; this result follows from the fact that for each data pair $(\bm{x}_{l},\bm{x}_{m})$ it is required $O(Dd)$ steps to find the angles $\{\sqrt{\gamma}\bm{\theta}_{j}\cdot\bm{x}\}$, $O(d)$ to prepare the QAFF states and $R$ measurement shots to estimate the projection. In Appendix B, we illustrate the overall quantum complexity of training QAFF by taking into account the gradients of the method.